# Equitability and explosive synchronisation in multiplex and higher-order networks

### v1

@@ V Independent cluster synchronisation implies equitability
    $\mathcal{P}$ is structurally equitable (Eq. (10)).
@@
(See the SI for a full derivation.) Since we have shown that dynamically equitability is a necessary condition for (linearly independent) cluster synchronisation, so is structural equitability. This restricts the groups of nodes in a network that can support cluster synchronisation to those forming equitable partitions simultaneously on each layer.
@@
Note that less restrictive notions of structural equitability are possible, for instance equitability with respect to all layers instead of per layer, or with respect to all external nodes instead of per cluster, but we can show that they are not necessary for cluster synchronisation (see SI for a full study on difference notions of equitability). Therefore, in the context of independent cluster synchronisation, a partition must be equitable on each layer simultaneously. This explains why cluster synchronisation is much harder to achieve in multi-layer networks: the equitability condition must be simultaneously satisfied on each layer, for the same nodes.
@@
Let us illustrate this simultaneous structural equitability condition with the two multilayer networks shown in Fig. 1. These are biplex networks, that is, multiplex networks (same nodes on all layers) with two layers. The first biplex network $G_{A}=(G_{A}^{(1)},G_{A}^{(2)})$ has a partition which is structurally (or externally) equitable in both layers simultaneously: the one with nodes $C=\{7,8,9,10\}$ as one cluster and all other nodes as singletons. Later, we show that cluster synchronisation does occur on $C$, on a numerical example (Fig. 2). On the other hand, the second biplex network $G_{B}=(G_{B}^{(1)},G_{B}^{(2)})$ violates the simultaneous equitable condition: even though there are almost identical external equitable partitions for every layer individually, $\{6,7,8,9,10,11\}$ on layer 1 and $\{7,8,9,10,11\}$ on layer 2, there are no partitions in which they are equitable for both layers (node 6 is essential to achieve equitability on layer 1, but prevents equitability on layer 2). Again, we show later on a numerical example (Fig.2) that only explosive synchronisation is observed in this case. This example illustrates how much harder it is to achieve equitability (and thus cluster synchronisation) in the multilayer setting.
@@
Figure 1: Two biplex networks $G_{A}=(G_{A}^{(1)},G_{A}^{(2)})$ on the left and $G_{B}=(G_{B}^{(1)},G_{B}^{(2)})$ on the right that exhibit drastically different dynamical behaviour: cluster synchronization on $G_{A}$, and explosive synchronization on $G_{B}$ (see Fig. 2 for the accompanying numerical results). This can be easily explained in terms of equitability: the green nodes form an equitable cluster in their respective layers, but only the common nodes $\{7,8,9,10\}$ are equitable overall on $G_{A}$, with no similar equitable partition on $G_{B}$.
@@
  • the partition of the underlying interaction network must be layer-by-layer equitable (Eq. (10)).
This summarises the key insight of our article: unless there are non-trivial linear dependencies between coupling functions (on trajectories), cluster synchronisation can only occur if the underlying partition is equitable.
In the next section, we show the converse statement, that is, that structural equitability (layer by layer) is not only necessary, but sufficient to guarantee the existence of a synchronised solution, with the help of the quotient network formalism.
Before moving on, we briefly remark on structural equitability and its relation to symmetry. Network symmetries (formally, permutation of the nodes preserving (weighted) edges) are an important source of equitable partitions, namely partitions into orbits of structurally equivalent nodes Sánchez-García (2020). For multilayer (multiplex) networks, and also for hypergraphs, we can extend the definition to permutations of the nodes preserving edges on each layer, respectively hyperedges of the same order. As an example, the equitable partition in the biplex $G_{A}$ in Fig. 1 is a symmetry partition into orbits.
VI Quotient dynamics
@@
Given the dynamical system (Eqs. (1) and (2)) and a partition $\mathcal{P}$ of the dynamical units (equivalently, of the set $\{1,\ldots,N\}$) into $K$ clusters, we define its quotient dynamical system with respect to $\mathcal{P}$ as the same dynamical system (same internal dynamics $\mathbf{f}$, coupling functions $\mathbf{g}^{(m)}$, and coupling strengths $\sigma_{m}$) on the quotient multilayer (multiplex) network, that is, the multiplex network with quotient adjacency matrices $B^{(m)}$, $m=1,\ldots,M$. Explicitly, we have dynamical units $\mathbf{y}_{1},\ldots,\mathbf{y}_{K}$ and $K$ differential equations
@@
This means that we can construct equitable, cluster synchronised solutions on an arbitrary multilayer network with respect to an arbitrary partition, simply by ‘lifting’ (repeating on each node in a cluster) a solution of the quotient multilayer network. Again, we are not making any claims on the stability of the synchronised solutions. Indeed, even if the quotient solution is stable, the parent solution will be stable under uniform perturbations of all the nodes in the cluster, but not necessarily under arbitrary, non-uniform perturbations.
@@
VII Numerical Results
We demonstrate the relation between equitability and cluster synchronisation with numerical simulations of coupled Lorenz oscillators on the two biplex network shown in Fig. 1 as we vary the coupling strength parameters $\sigma_{1}$ and $\sigma_{2}$. Full details of the simulations, including the dynamical functions involved, which belong to Class II Boccaletti et al. (2006) on each layer (global synchronisation guaranteed for high enough values of $\sigma_{i}$), and the definition of the synchronisation error, can be found in the SI.
@@ VII Numerical Results
The general results in our paper fully explain the numerically observed behaviour: the nodes $\{7,8,9,10\}$ form a layer-by-layer equitable partition on $G_{A}$, but no such partition exists on $G_{B}$ (Fig. 1), exemplifying the prevalence of explosive synchronisation on multilayer (and higher-order) networks compared to standard (monolayer) networks.
Figure 2: Synchronization error of coupled Lorenz oscillators in the two biplex networks shown in Fig. 1 with respect to the cluster $\{7,8,9,10\}$. Panels A1 and A2 (respectively B1 and B2) show the cluster synchronization in $G_{A}$ (respectively $G_{B}$) when only layer 2, or layer 1, is active. Panels A3 and B3 show the $10^{-5}$ contours of the synchronization errors for $(\sigma_{1},\sigma_{2})\in[0,10]\times[0,10]$ for networks $G_{A}$ and $G_{B}$, respectively. See the SI for full details.
VIII Results for hypergraph dynamics
@@
We also define corresponding dynamical quantities $\mathbf{h}_{i}^{C_{1},\ldots,C_{m}}$, given a solution $\vec{\mathbf{x}}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})$, by adding $\mathbf{g}^{(m)}(\mathbf{x}_{i},\mathbf{x}_{j_{1}},\ldots,\mathbf{x}_{j_{m}})$ to Eq. (18) above. These account for the total structural, respectively dynamical, input to node $i$ from all nodes in $C_{1},\ldots,C_{m}$.
@@
where $\mathbf{h}_{i}^{\text{ext}}=\mathbf{h}_{i}-\sum_{m=1}^{M}\sigma_{m}\,\mathbf{h}_{i}^{C,\stackrel{{\scriptstyle m}}{{\ldots}},C}$ with $C,\stackrel{{\scriptstyle m}}{{\ldots}},C$ representing $C$ repeated $m$ times. We call the partition $\mathcal{P}$ structurally equitable if, for all $C\in\mathcal{P}$, $i,j\in C$, we have
@@
As in the multilayer case, we can show that cluster synchronisation implies dynamical equitability, and an equivalence between dynamical and structural equitability for independent cluster synchronisation (see the SI).
This shows why we expect explosive synchronisation to be prevalent in generic higher-order dynamics as well: for a partition to be structurally equitable (a necessary condition for independent cluster synchronisation), the equitability condition must be satisfied simultaneously for each subset of $m$ clusters for all $m$ (Eq. (21)).
To demonstrate our theoretical results numerically, we simulated Lorenz oscillators in a hypergraph with 20 nodes, four clusters, and both 2- and 3-body interactions (see the SI for a full description of the system). The clusters were chosen such that $C_{1}$ is (externally) equitable with respect to pairwise but not triadic interactions; cluster $C_{2}$ is equitable for triadic but not pairwise interactions; cluster $C_{3}$ satisfy neither pairwise nor triadic equitability; and cluster $C_{4}$ is the only one that simultaneously satisfy pairwise and triadic equitability. The results are shown on Fig. 3, which shows the mean square error of the four clusters for a fixed and small value of the triadic interaction coupling ($\sigma_{2}=0.2$), and varying pairwise interaction coupling ($\sigma_{1}$). We can see that only cluster $C_{4}$ synchronises before the whole system does. We also note that cluster $C_{1}$ becomes almost synchronous due to triadic interactions becoming small with respect to the pairwise interactions for which $C_{1}$ does have structural equitability.
Figure 3: Synchronization error of coupled Lorenz oscillators for four cluster on a hypergraph (fully described in the SI) with 20 nodes and both pairwise and triadic interactions. The clusters satisfy only pairwise respectively triadic structural equitability (clusters $C_{1}$ respectively $C_{2}$), none (cluster $C_{3}$) or simultaneously both (cluster $C_{4}$). The coupling parameter $\sigma_{2}$ is fixed at 0.2 while $\sigma_{1}$ is increased. Only cluster $C_{4}$ synchronises before the whole system does, as shown by the synchronisation error, which agrees with our prediction, as the only simultaneously equitable cluster for all types of interactions. Cluster $C_{1}$ becomes almost synchronous due to the triadic interactions becoming small with respect to the pairwise interactions for which $C_{1}$ does have structural equitability.
Finally, we can also prove the reciprocal statement, that is, that the structural equitability of a given partition implies the existence of cluster synchronised solutions, by ‘lifting’ (repeating on each cluster) a solution of the quotient dynamics to the parent hypergraph. For completeness, we give below the definitions of quotient dynamics for multi-body interactions, with full details included in the SI.
@@
IX Conclusions

### v3

@@ II.4 Independent cluster synchronisation implies external equitability
    $\mathcal{P}$ is externally equitable (Eq. (10)).
@@
(See the SI for a full derivation.) Since we have shown that dynamically equitability is a necessary condition for (linearly independent) cluster synchronisation, so is external equitability. This restricts the groups of nodes in a network that can support cluster synchronisation to those forming equitable partitions simultaneously on each layer. (In the next section, we will see that external equitability is also a sufficient condition for cluster synchronisation.)
@@
We could have considered weaker, less restrictive, notions of external equitability, for instance equitability with respect to all layers instead of per layer, or with respect to all external nodes instead of per cluster, but we can show that they are not necessary for cluster synchronisation (see SI for a full comparison on difference notions of equitability). Therefore, in the context of independent cluster synchronisation, a partition must be equitable on each layer simultaneously. This explains why cluster synchronisation is much harder to achieve in multi-layer networks: the equitability condition must be simultaneously satisfied on each layer, for the same nodes. Indeed, even if many networks have equitable partitions with high probability somewhere on the network (for example, orbit partitions in real-world networks macarthur2008symmetry ; sanchez2020exploiting ), the probability of equitable clusters to occur on the same nodes in two, or more, layers decays very rapidly with the number of layers (see the SI for an explicit calculation).
@@
Let us illustrate the simultaneous external equitability condition with the two multilayer networks shown in Fig. 2. These are biplex networks, that is, multiplex networks (same nodes on all layers) with two layers. The first biplex network $G_{A}=(G_{A}^{(1)},G_{A}^{(2)})$ has a partition which is externally equitable in both layers simultaneously: the one with nodes $C=\{7,8,9,10\}$ as one cluster and all other nodes as singletons. Later, we show that cluster synchronisation does occur on $C$, on a numerical example (Fig. 3). On the other hand, the second biplex network $G_{B}=(G_{B}^{(1)},G_{B}^{(2)})$ violates the simultaneous equitable condition: even though there are almost identical external equitable partitions for every layer individually, $\{6,7,8,9,10,11\}$ on layer 1 and $\{7,8,9,10,11\}$ on layer 2, there are no partitions which are equitable on both layers simultaneously (node 6 is essential to achieve equitability on layer 1, but prevents equitability on layer 2). Again, we show later on a numerical example (Fig. 3) that only explosive synchronisation is observed in this case.
@@
Figure 2: Equitability on multiplex networks. Two biplex networks $G_{A}=(G_{A}^{(1)},G_{A}^{(2)})$ on the left and $G_{B}=(G_{B}^{(1)},G_{B}^{(2)})$ on the right that exhibit very different dynamical behaviour: cluster synchronization on $G_{A}$, and explosive synchronization on $G_{B}$ (see Fig. 3 for the accompanying numerical results). This can be easily explained in terms of equitability: the green nodes form an equitable cluster in their respective layers, but only the common nodes $\{7,8,9,10\}$ are equitable overall on $G_{A}$, with no similar equitable partition on $G_{B}$.
@@
  • the partition of the underlying interaction network must be layer-by-layer externally equitable (Eq. (10)).
This summarises the key insight of our article: unless there are linear dependencies between coupling functions (on trajectories), cluster synchronisation can only occur if the underlying partition is externally equitable layer by layer. (For a comparison to other possible notions of equitability, see the SI.)
In the next section, we show the converse statement, that is, that external equitability (layer by layer) is not only necessary, but sufficient to guarantee the existence of a synchronised solution, with the help of the quotient network formalism.
Before moving on, we briefly remark on external equitability and its relation to symmetry. Network symmetries (formally, permutation of the nodes preserving (weighted) edges) are an important source of equitable partitions, namely partitions into orbits of structurally equivalent nodes macarthur2008symmetry ; sanchez2020exploiting . For multiplex networks, and also for hypergraphs, we can extend the definition to permutations of the nodes preserving edges on each layer, respectively hyperedges of the same order. As an example, the equitable partition in the biplex $G_{A}$ in Fig. 2 is indeed a symmetry partition into orbits.
II.5 Quotient dynamics
Quotients with respect to equitable partitions have been studied in the context of cluster synchronisation and dynamics e.g. golubitsky2005patterns ; sorrentino2019symmetries ; aguiar2023network . Our treatment of quotient and quotient dynamics here is similar or equivalent to that found elsewhere.
@@ II.5 Quotient dynamics
Given the dynamical system (Eqs. (1) and (2)) and a partition $\mathcal{P}$ of the dynamical units (equivalently, of the set $\{1,\ldots,N\}$) into $K$ clusters, we define its quotient dynamical system with respect to $\mathcal{P}$ as the same dynamical system (same internal dynamics $\mathbf{f}$, coupling functions $\mathbf{g}^{(m)}$, and coupling strengths $\sigma_{m}$) on the quotient multiplex network, that is, the multiplex network with quotient adjacency matrices $B^{(m)}$, $m=1,\ldots,M$. Explicitly, we have dynamical units $\mathbf{y}_{1},\ldots,\mathbf{y}_{K}$ and $K$ differential equations
@@
This means that we can construct equitable, cluster synchronised solutions on an arbitrary multiplex network with respect to an arbitrary partition, simply by ‘lifting’ (repeating on each node in a cluster) a solution of the quotient multiplex network. That is, in this sense, equitability implies cluster synchronisation (the existence of cluster synchronised solutions on the clusters of the chosen equitable partition). Again, we are not making any claims on the stability of the synchronised solutions: even if the quotient solution is stable, the parent solution will be stable under uniform perturbations of all the nodes in the cluster, but not necessarily under arbitrary, non-uniform perturbations.
@@
II.6 Numerical Results
We illustrate the relation between equitability and cluster synchronisation with numerical simulations of coupled Lorenz oscillators on the two biplex network shown in Fig. 2 as we vary the coupling strength parameters $\sigma_{1}$ and $\sigma_{2}$. The quantity of interest we track here is the synchronization error,
$S_{k}=\left\langle\sqrt{\frac{1}{N_{k}}\sum_{i\in C_{k}}\left|\mathbf{x}_{i}-\overline{\mathbf{x}}_{k}\right|^{2}}\right\rangle_{T},$ (18)
where $C_{k}$ is the set of nodes contained in cluster $k$ (or all nodes in the case of the total error), $\overline{\mathbf{x}}_{k}$ is the ensemble average of $\vec{\mathbf{x}}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})$ within $C_{k}$, $\overline{\mathbf{x}}_{k}=\frac{1}{N_{k}}\sum_{i\in C_{k}}\mathbf{x}_{i}$, where $N_{k}=|C_{k}|$ the number of nodes in $C_{k}$, and $\langle\cdot\rangle_{T}$ denotes the temporal average over a late time window $T$. Full details of the simulations, including the dynamical functions involved, which belong to Class II boccaletti2006structure on each layer (global synchronisation guaranteed for high enough values of $\sigma_{i}$), can be found in the SI.
@@ II.6 Numerical Results
The general results in our paper fully explain the numerically observed behaviour: the nodes $\{7,8,9,10\}$ form a layer-by-layer equitable partition on $G_{A}$, but no such partition exists on $G_{B}$ (Fig. 2), exemplifying the prevalence of explosive synchronisation on multilayer (and higher-order) networks compared to standard (monolayer) networks. (See also the ‘Explosive synchronisation and equitability’ section below.)
Figure 3: Equitability and synchronisation on two multiplex networks. Synchronization error Eq. (18) of coupled Lorenz oscillators in the two biplex networks shown in Fig. 2 with respect to the cluster $\{7,8,9,10\}$. Panels A1 and A2 (respectively B1 and B2) show the cluster synchronization errors in $G_{A}$ (respectively $G_{B}$) when only layer 2, or layer 1, is active. Panels A3 and B3 show the $10^{-5}$ contours of the synchronization errors (i.e. the regions whose error is below or above $10^{-5}$) for $(\sigma_{1},\sigma_{2})\in[0,10]\times[0,10]$ for networks $G_{A}$ and $G_{B}$, respectively. See the SI for full details.
II.7 Results for hypergraph dynamics
@@
$h_{i}^{C_{1},\ldots,C_{m}}$ $=\sum_{j_{1}\in C_{1}}\ldots\sum_{j_{1}\in C_{m}}A^{(m)}_{i,j_{1},\ldots,j_{m}}$. (19)
We also define corresponding dynamical quantities $\mathbf{h}_{i}^{C_{1},\ldots,C_{m}}$, given a solution $\vec{\mathbf{x}}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})$, by adding $\mathbf{g}^{(m)}(\mathbf{x}_{i},\mathbf{x}_{j_{1}},\ldots,\mathbf{x}_{j_{m}})$ to Eq. (19) above. These quantities account for the total structural, respectively dynamical, input to node $i$ from all nodes in $C_{1},\ldots,C_{m}$.
@@
where $\mathbf{h}_{i}^{\text{ext}}=\mathbf{h}_{i}-\mathbf{h}_{i}^{\text{int}}=\mathbf{h}_{i}-\sum_{m=1}^{M}\sigma_{m}\,\mathbf{h}_{i}^{C,\stackrel{{\scriptstyle m}}{{\ldots}},C}$ with $C,\stackrel{{\scriptstyle m}}{{\ldots}},C$ representing $C$ repeated $m$ times. We call the partition $\mathcal{P}$ externally equitable if, for all $C\in\mathcal{P}$, $i,j\in C$, we have
@@
As in the multiplex case, we can show that cluster synchronisation implies dynamical equitability, and an equivalence between dynamical and external equitability for independent cluster synchronisation (see the SI for precise mathematical statements and proofs). In particular, we have that cluster synchronisation can only occur in one of these cases:
  • there is a linear dependence relation between the coupling functions on trajectories; or, if not,
  • the partition of the underlying interaction hypergraph must be externally equitable (Eq. (21)).
To demonstrate our theoretical results numerically, we simulated Lorenz oscillators in a hypergraph with 20 nodes, four clusters, and both 2- and 3-body interactions (see the SI for a full description of the system). The clusters were chosen such that $C_{1}$ is (externally) equitable with respect to pairwise but not triadic interactions; cluster $C_{2}$ is equitable for triadic but not pairwise interactions; cluster $C_{3}$ satisfy neither pairwise nor triadic equitability; and cluster $C_{4}$ is the only one that simultaneously satisfy pairwise and triadic equitability. The results (Fig. 4) show the mean square error of the four clusters for a fixed and small value of the triadic interaction coupling ($\sigma_{2}=0.2$), and varying pairwise interaction coupling ($\sigma_{1}$). We can see that only cluster $C_{4}$ synchronises before the whole system does. We also note that cluster $C_{1}$ becomes almost synchronous due to triadic interactions becoming small with respect to the pairwise interactions for which $C_{1}$ does satisfy external equitability.
Figure 4: Equitability and synchronisation on a hypergraph. Synchronization error of coupled Lorenz oscillators for four cluster on a hypergraph (fully described in the SI) with 20 nodes and both pairwise and triadic interactions. The clusters satisfy only pairwise respectively triadic structural equitability (clusters $C_{1}$ respectively $C_{2}$), none (cluster $C_{3}$) or simultaneously both (cluster $C_{4}$). The coupling parameter $\sigma_{2}$ is fixed at 0.2 while $\sigma_{1}$ is increased. Only cluster $C_{4}$ synchronises before the whole system does, as shown by the synchronisation error, which agrees with our prediction, as the only simultaneously equitable cluster for all types of interactions. Cluster $C_{1}$ becomes almost synchronous due to the triadic interactions becoming small with respect to the pairwise interactions for which $C_{1}$ does have external equitability.
Finally, we can also prove the reciprocal statement, that is, that the external equitability of a given partition implies the existence of cluster synchronised solutions, by ‘lifting’ (repeating on each cluster) a solution of the quotient dynamics to the parent hypergraph. For completeness, we give below the definitions of quotient dynamics for multi-body interactions, with full details included in the SI.
@@
Similarly to the multiplex case, we show (in the SI) that $(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})$ is a $\mathcal{P}$-synchronised solution of the parent dynamics if and only if $(\mathbf{y}_{1},\ldots,\mathbf{y}_{K})$ is a solution of the quotient dynamics, where $\mathbf{y}_{k}=\mathbf{x}_{i}$ whenever $i\in C_{k}\in\mathcal{P}$. This means that any solution of the quotient system (with respect to an externally equitable partition) gives rise to a cluster synchronised solution (with respect to the clusters in the partition) of the original (parent) system.
II.8 General coupling and general equitability
Our approach can be adapted to general coupling functions, that is, dropping the non-invasiveness condition $\mathbf{g}(x,\ldots,x)=0$ in Eqs. (2) and (3), by substituting external equitability with general equitability. That is, we need to consider the internal cluster connectivity which is not necessarily zero when the cluster is synchronised in this general case.
Our previous results can be extended to general coupling functions with the following changes from external to general equitability:
  • substitute (8) and (20) by $\mathbf{h}_{i}=\mathbf{h}_{j}$;
  • remove $C\neq C^{\prime}$ before (10) and (11);
  • remove ‘not all equal to $C$’ after (21) and (22).
With these changes, our results hold: for linear independent solutions, cluster synchronisation implies (general) equitability and vice-versa (see the SI for precise mathematical statements and derivations).
Perhaps counter-intuitively at first, the generic case includes the non-invasive case but needs to be treated separately: if any of the coupling functions are non-invasive (any of the functions satisfy $\mathbf{g}(\mathbf{x},\ldots,\mathbf{x})=0$, the solutions are never linearly independent (the family of functions Eq. (11) includes a zero function when $C=C^{\prime}$) and our results do not apply (no linearly independent solution exists). If all functions are non-invasive (they all satisfy $\mathbf{g}(\mathbf{x},\ldots,\mathbf{x})=0$), our previous approach (first part of the paper) applies, by restricting to external equitability.
More generally, if we have a mixed system with non-invasive and general coupling functions, our results apply to the corresponding mixed notion of equitability: external equitability on layer $m$, respectively order $m$, if $g^{(m)}$ is non-invasive, otherwise general equitability, and adapting the notion of linear independence condition (remove $\mathbf{g}^{(m)}(\mathbf{x}_{C},\mathbf{x}_{C})$ whenever $\mathbf{g}^{(m)}$ is non-invasive). The same also applies to a general coupling function that becomes non-invasive on a particular synchronised trajectory ($\mathbf{g}(\mathbf{x}_{C},\mathbf{x}_{C})=0$ for some cluster $C$).
II.9 Explosive synchronisation and equitability
Explosive synchronisation refers to the sudden transition from a non-synchronised to a fully synchronised state of the dynamical units. Our results guarantee that explosive synchronisation is the only possible generic (that is, linearly independent) synchronisation that can occur if the only equitable partitions are partitions into 1 or $N$ clusters (corresponding to global, respectively no synchronisation). Let us refer to this as generic explosive synchronisation, that is, systems for which the only generic (linearly independent) solution is the globally synchronised one ($\mathbf{x}_{i}=\mathbf{x}_{j}$ for all $i,j$).
In the non-invasive case, the partition into one cluster is always externally equitable (condition (10) is empty, as there is no external connectivity to consider) and, indeed, any solution of the internal dynamics $\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})$ is a globally synchronised solution of the system, as explained earlier. Therefore, we have generic explosive synchronisation if and only if the underlying partition have no external equitable partitions except the partitions into one and into $N$ clusters — a purely structural condition of the underlying interaction multiplex or hypergraph.
In the general case, however, the partition into one cluster may not be equitable, as we now need to consider internal equitability. Indeed, the system admits a globally synchronised solution if and only if the partition into one cluster is equitable. This, in turn, is equivalent to a regularity condition on the underlying multiplex, or hypergraph, namely that all nodes have the same weighted in-degree on each layer, respectively for each order $m$ (see the SI for full details).
II.10 Finding equitable partitions
Finding all equitable partition on an arbitrary graph (network) is in general a computationally hard problem lubiw1981some , but there are some tractable cases. For example, the refinement procedure in the Weisfeiler-Lehmann (WL) algorithm huang2021short finds the coarsest equitable partition in polynomial time — any other equitable partition must be a refinement (obtained but further subpartitioning the clusters). Any other equitable partition must therefore be a refinement of the WL one. In particular, only nodes within the same WL cluster can synchronise (in a generic, that is, linearly independent, way). The WL refinement algorithm can be easily adapted to weighted networks, multiplexes, and hypergraphs, and to obtain the coarsest equitable refinement of a given partition $\mathcal{P}$ (see SI for details).
The case of external equitability is more subtle. Since the computation of all equitable partitions is computationally hard, so is the computation of all external equitable partitions. Any equitable partition, including the WL one, is also externally equitable, but there can be more (for example, the partition into one cluster is always externally equitable). The WL algorithm can be adapted to find the coarsest externally equitable partition containing a given partition $\mathcal{P}$ (see the SI) but it cannot be used to find the coarsest partition overall: the partition into one cluster (whole network) is already externally equitable.
It would be very interesting to find algorithms to find equitable and externally equitable partitions in the large but sparse graphs typically found in applications, but this is out of the scope of the present article.
III Conclusions